# Fielding: Clustered Federated Learning with Data Drift

### v2

@@ Appendix C Implementation
We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine. We follow FedScale’s design of having one centralized coordinator and client-specific executors. Figure 11 illustrates the end-to-end workflow of the Fielding system prototype with label distribution vectors as client representations. At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally. The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts. After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership. In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987). The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
@@ E.1 Gradients as representation
As mentioned in Section 2, we propose using label distribution vectors as client representations when we deal with label and covariate shifts. Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes. Compared to gradient-based clustering, label-based doesn’t demand parameter or gradient transmission, doesn’t introduce additional computation tasks, and is not sensitive to the timing of client clustering. In Table 3, label-based clustering provides consistent heterogeneity reduction across rounds. Furthermore, as a label distribution vector typically has tens or hundreds of coordinates compared to millions in a full gradient, label-based clustering incurs significantly lower computational overhead. In our experiment of training ResNet-18 on the FMoW dataset, clustering all 302 clients using our label-based solution takes only 0.05 seconds while the gradient-based solution FlexCFL takes 37.92 seconds (note that FlexCFL already accelerates this process through dimensionality reduction using truncated Singular Value Decomposition).
@@
In terms of overhead, both embedding- and gradient-based clustering introduce client-side computation, which can be mitigated by deploying a smaller shared model for representation collection, eliminating extra model download time. We adopt this approach when we gather the time-to-accuracy results of training ResNet-18 on Cityscapes with various representations (shown in Figure 13). We first construct a small training set by randomly sampling 200 images from each class. We then train a ResNet-18 model on this dataset for 300 epochs and broadcasts it so that clients store it locally and use it for gradient and embedding generation. Note that gradient-based Fielding still takes longer to finish 200 training rounds than other variants in Figure 13 due to the longer computational time of back propagation and longer transmission time of gradient vectors.
Handling concept drifts where $P(y|x)$ changes requires loss-based representations, making gradients more appropriate than label distribution vectors or embeddings. Figure 13 shows the time-to-accuracy results with synthesized concept drifts on Cityscapes. Here we make all data samples available throughout the training, but introduce drift events by randomly choosing 50% clients and having each chosen client randomly pick two labels and swap their samples (i.e., if a client picks label $A$ and $B$, then all samples previously labeled $A$ are now labeled $B$, and samples labeled $B$ now have label $A$). Gradient-based Fielding manages to retain test accuracy under such aggressive concept drift, while label and embedding-based Fielding have similar results as the baseline without any clustering. This result highlights that gradient has the potential of addressing concept drifts while label distribution and embedding don’t.
@@ E.2 Input embeddings as representations
Since covariate shift is defined as input distribution ($P(x)$) changing while the feature-to-label mapping ($P(y|x)$) remains constant, input embedding distance should be a proxy for label distribution distance. Hence, minimizing the distance between input embeddings naturally minimizes the label distribution distance as well. Table 3 supports this intuition by showing a correlation between the label distribution distance and the embedding distance. This observation suggests that using label distribution vectors or input embeddings as representations should have comparable performance on labeled datasets. As shown in Figure 13, both label distribution vectors and embeddings enable Fielding to outperform the no-clustering baseline (the sudden accuracy drop of the embedding-based curve around 150 minutes is the result of global re-clustering). Input embeddings have the potential of capturing label and covariate shifts on unlabeled data, which label distribution vectors are incapable of.
@@ F.2 Alternative global clustering threshold using pairwise distance
To demonstrate that Fielding prototype performs empirically well with the alternate re-clustering thresholds presented in Appendix A, we re-run the Cityscapes task with re-clustering mechanism matching Line 9-11 of Algorithm 4. In specific, we measure clients’ pairwise distance as the L1 distance between their distribution vectors, and tune $\Delta$ adaptively. We start with $\Delta=c=0.1$. After each data drift event, we check whether global re-clustering has been triggered consecutive by two drift events. If so, we update $\Delta$ to $2\times\Delta$; otherwise, we update $\Delta$ to $\min(c,\Delta-c)$. Figure 16 shows that the prototype of the Fielding system works well with both center shift distance-based and client pairwise distance-based global re-clustering triggering conditions.
Figure 15: Fielding performance with different global re-clustering triggering conditions (center shift distance as in Algorithm 2 or client pairwise distance as in Appendix A).
@@ F.3 Supporting diverse distance metric
To demonstrate that Fielding is not tied to specific distance metric, we run the Cityscapes task with label distribution vectors as client representations, and both L1 and Jensen–Shannon distance (Lin, 1991) as the distance metric. Figure 16 shows that Fielding is compatible with different distance metrics chosen for a given client representation.

### v3

@@ Appendix C Implementation
We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine. We follow FedScale’s design of having one centralized coordinator and client-specific executors. Figure 11 illustrates the end-to-end workflow of the Fielding system prototype with label distribution vectors as client representations. At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally. The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts. After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership. In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987). The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm˜2).
@@ E.1 Gradients as representation
As mentioned in Section˜2, we propose using label distribution vectors as client representations when we deal with label and covariate shifts. Label distribution vectors are available on all clients and enable us to promptly detect any drift by checking for distribution changes. Compared to gradient-based clustering, label-based doesn’t demand parameter or gradient transmission, doesn’t introduce additional computation tasks, and is not sensitive to the timing of client clustering. In Table 3, label-based clustering provides consistent heterogeneity reduction across rounds. Furthermore, as a label distribution vector typically has tens or hundreds of coordinates compared to millions in a full gradient, label-based clustering incurs significantly lower computational overhead. In our experiment of training ResNet-18 on the FMoW dataset, clustering all 302 clients using our label-based solution takes only 0.05 seconds while the gradient-based solution FlexCFL takes 37.92 seconds (note that FlexCFL already accelerates this process through dimensionality reduction using truncated Singular Value Decomposition).
@@
In terms of overhead, both embedding- and gradient-based clustering introduce client-side computation, which can be mitigated by deploying a smaller shared model for representation collection, eliminating extra model download time. We adopt this approach when we gather the time-to-accuracy results of training ResNet-18 on Cityscapes with various representations (shown in Figure˜13). We first construct a small training set by randomly sampling 200 images from each class. We then train a ResNet-18 model on this dataset for 300 epochs and broadcasts it so that clients store it locally and use it for gradient and embedding generation. Note that gradient-based Fielding still takes longer to finish 200 training rounds than other variants in Figure˜13 due to the longer computational time of back propagation and longer transmission time of gradient vectors.
Handling concept drifts where $P(y|x)$ changes requires loss-based representations, making gradients more appropriate than label distribution vectors or embeddings. Figure˜13 shows the time-to-accuracy results with synthesized concept drifts on Cityscapes. Here we make all data samples available throughout the training, but introduce drift events by randomly choosing 50% clients and having each chosen client randomly pick two labels and swap their samples (i.e., if a client picks label $A$ and $B$, then all samples previously labeled $A$ are now labeled $B$, and samples labeled $B$ now have label $A$). Gradient-based Fielding manages to retain test accuracy under such aggressive concept drift, while label and embedding-based Fielding have similar results as the baseline without any clustering. This result highlights that gradient has the potential of addressing concept drifts while label distribution and embedding don’t.
@@ E.2 Input embeddings as representations
Since covariate shift is defined as input distribution ($P(x)$) changing while the feature-to-label mapping ($P(y|x)$) remains constant, input embedding distance should be a proxy for label distribution distance. Hence, minimizing the distance between input embeddings naturally minimizes the label distribution distance as well. Table 3 supports this intuition by showing a correlation between the label distribution distance and the embedding distance. This observation suggests that using label distribution vectors or input embeddings as representations should have comparable performance on labeled datasets. As shown in Figure˜13, both label distribution vectors and embeddings enable Fielding to outperform the no-clustering baseline (the sudden accuracy drop of the embedding-based curve around 150 minutes is the result of global re-clustering). Input embeddings have the potential of capturing label and covariate shifts on unlabeled data, which label distribution vectors are incapable of.
@@ F.2 Alternative global clustering threshold using pairwise distance
To demonstrate that Fielding prototype performs empirically well with the alternate re-clustering thresholds presented in Appendix˜A, we re-run the Cityscapes task with re-clustering mechanism matching Line 9-11 of Algorithm˜4. In specific, we measure clients’ pairwise distance as the L1 distance between their distribution vectors, and tune $\Delta$ adaptively. We start with $\Delta=c=0.1$. After each data drift event, we check whether global re-clustering has been triggered consecutive by two drift events. If so, we update $\Delta$ to $2\times\Delta$; otherwise, we update $\Delta$ to $\min(c,\Delta-c)$. Figure˜16 shows that the prototype of the Fielding system works well with both center shift distance-based and client pairwise distance-based global re-clustering triggering conditions.
Figure 15: Fielding performance with different global re-clustering triggering conditions (center shift distance as in Algorithm˜2 or client pairwise distance as in Appendix˜A).
@@ F.3 Supporting diverse distance metric
To demonstrate that Fielding is not tied to specific distance metric, we run the Cityscapes task with label distribution vectors as client representations, and both L1 and Jensen–Shannon distance (Lin, 1991) as the distance metric. Figure˜16 shows that Fielding is compatible with different distance metrics chosen for a given client representation.